OpenStudy (anonymous):

let f(x) = 80 + 4x − 4x2 and g(x) = 4x2 − 20x. Sketch the region enclosed by the graphs of f(x) and g(x) and compute its area.

OpenStudy (anonymous):

did you graph them yet? when you do, locate their points of intersection. these will be your limits of integration

OpenStudy (anonymous):

i got 457.333 but it's not right.

OpenStudy (anonymous):

What did you get for the points of intersection ?

OpenStudy (anonymous):

i took 80+4x-4x^2=4x^2-20x and then equalled it to zero

OpenStudy (anonymous):

0=8x^2-24x-80

OpenStudy (anonymous):

0=8(x^2-3x-10) 0=8(x+2)(x-5) x=-2,5

OpenStudy (anonymous):

perfect, these will be the limits of integration next we just determine which function is larger on that interval

OpenStudy (anonymous):

\[\int\limits_{-2}^{5}(80+24x-8x^2)dx\]

OpenStudy (anonymous):

what do i do now though?

OpenStudy (anonymous):

integrate it it becomes 80x + 12x^2 - 8/3 x^3

OpenStudy (anonymous):

(80(5)+12(5)^2-8/3(5)^3)-(80(-2)+12(-2)^2-8/3(-2)^3)=457.333

OpenStudy (anonymous):

thats it. maybe if you dont want to estimate just use 1372/3

OpenStudy (anonymous):

i just entered it onto my online homework and that was the problem. thank you so much.

OpenStudy (anonymous):

yeah those things are picky :\

OpenStudy (anonymous):

looks like they are midpoints, each with a base of 1 so calculate the function's value at those midpoints and determine the area that way

OpenStudy (anonymous):

just plug the -1,-2,-3,1,2,3 into the x in the equation and add them up?

OpenStudy (anonymous):

well no the base for all rectangles is 1 but their heights are the midpoints evaluated at the function

OpenStudy (anonymous):

so your x values would be -2.5, -1.5, -0.5, 0.5, 1.5, 2.5